A progressive wave is given T=0.52π(40t-20x) find



a)amplitude



b)frequency



c)wavespeed



e)wavelenght

1
Expert's answer
2022-03-27T18:20:29-0400

"A=0.52\pi,"

"\omega=40,"

"v=A\omega=20.8\pi,"

"\lambda=80\pi."